OpenStudy (anonymous):

rationalize the denominator 3+√3/4-√3

OpenStudy (anonymous):

multiply the top and bottom by the conjugate of the denominator, which is 4+sqrt3 \[\frac{3+\sqrt{3}}{4-\sqrt{3}}*\frac{4+\sqrt{3}}{4+\sqrt{3}}=\frac{15+7\sqrt{3}}{13}\] (some "foiling" omitted)

OpenStudy (anonymous):

it is right but where did you get the 13 from??

OpenStudy (anonymous):

ok, that was part of the omitted foiling, which I'll now show. in the bottom we have \[(4-\sqrt{3})(4+\sqrt{3})=16+4\sqrt{3}-4\sqrt{3}-3=16-3=13\]

OpenStudy (anonymous):

on the roots cancelled out ...alrighty